Find the rates of convergence of the following functions as h → 0.
lim
h→0
sin (h − h cos h) /h
= 0
1
Expert's answer
2019-03-21T10:08:32-0400
"\lim_{h\to0} \frac{sin(h-hcos(h))}{h}"

So, we know that


"\lim_{h\to0} \frac{sin(h)}{h}=1"

Let's multiply the numerator and denominator by (h-hcos(h)):


"\lim_{h\to0} \frac{sin(h-hcos(h))*(h-hcos(h))}{h*(h-hcos(h))} = (1)""\lim_{h\to0}\frac{sin(h-hcos(h))}{h-hcos(h)}=1"

"(1) = \lim_{h\to0} \frac{h-hcos(h)}{h}=\lim_{h\to0}1-cos(h)"

Use the Maclaurin series for cosine:


"cos(h) = 1-\frac{h^{2}}{2!}+\frac{h^{4}}{4!}-\frac{h^{6}}{6!}+..."

"1-cos(h)=1-(1-\frac{h^{2}}{2!}+...)=\frac{h^{2}}{2}+...=O(h^{2})"

Answer: the function has quadratic convergence.